# Unsupervised learning in echo state networks for input reconstruction

### v3

@@ 3.2.3 Numerical experiments
We generated inputs $d_{t}^{(1)}=d_{t}+\epsilon_{t}^{(1)}$ and $d_{t}^{(2)}+\epsilon_{t}^{(1)}$, where $\epsilon_{t}^{(1)}$ and $\epsilon_{t}^{(2)}$ are zero-mean white Gaussian noise with variance $\text{tr}(\Sigma^{(1)})=0.01$ and $\text{tr}(\Sigma^{(1)})\in$ numpy.logspace(-2, 1, 10):10 numbers spaced evenly on a log scale between $10^{-2}$ and $10^{1}$. We used a 10-dimensional linear ESN with the identity function $\text{id}:r\mapsto r$ as the activation function $\sigma$. Let $A_{0}\in\mathbb{R}^{10\times 10}$ be a matrix with elements being i.i.d samples from $\mathcal{N}(0,1)$. We set $A=0.9\cdot\lambda(A_{0})\cdot A_{0}$, where $\lambda(A_{0})$ is the spectral radius of $A_{0}$. We set the elements of the input layer parameter $B\in\mathbb{R}^{10\times 1}$ using i.i.d samples obtained from $\mathcal{N}(0,0.02)$. For each ESN implementation and noisy inputs $\mathcal{D}_{1,2000}^{(1)}$ and $\mathcal{D}_{1,2000}^{(2)}$, we calculated reservoir states $\mathcal{R}_{1,2001}^{(1)}$ and $\mathcal{R}_{1,2001}^{(2)}$ with the zero vector as the initial reservoir state. We applied Algorithm 2 to obtain $W_{R}^{(1)},\hat{f}_{\mathcal{R}}^{(1)}$ and $\hat{Q}$ from $\mathcal{R}_{1,2000}^{(1)}$. Then, we applied the adaptive Kalman filter Algorithm 3 with $\alpha_{R}=0.1$ to obtain filtered reservoir states $r_{t}^{(2\to 1)}$ from the sequence of contaminated reservoir states $r_{t}^{(2)}$. We conducted simulations for 50 different implementations of the ESN parameters $A$ and $B$ and noisy inputs $\mathcal{D}_{1,2000}^{(1)}$ and $\mathcal{D}_{1,2000}^{(2)}$. We calculated the relative root mean squared error (RRMSE) to evaluate the IR performance.
@@
Figure 6 shows a typical numerical example of the ESN inputs and outputs using IR with noise filtering. We used low-intensity (Figure 6(a)) and high-intensity (Figure 6(b)) noise inputs for training and testing, respectively. In this trial, the RRMSE of IR was reduced from $1.41$ to $0.57$ by applying noise filtering (Figure 6(c) and (d)).
@@
Figure 6: Typical example of input reconstruction with noise filtering. a. Time series of the ESN input $d_{t}^{(1)}$ with low-intensity noise $(\text{tr}(\Sigma^{(1)})=0.01)$ used during prior training b. Time series of the ESN input $d_{t}^{(2)}$ with high-intensity noise $(\text{tr}(\Sigma^{(2)})=1)$ used during testing c. Time series of the ESN output $W_{\mathcal{R}}^{(1)}r_{t}^{(2)}$ before filtering d. Time series of the ESN output $W_{\mathcal{R}}^{(1)}r_{t}^{(2\to 1)}$ after filtering. For all plots in ad, the time series of the original ESN input $d_{t}$ is plotted for reference (black dashed line).
@@
Figure 7 shows the average behavior of the IR RRMSE across 50 trials for each noise intensity of inputs $d_{t}^{(2)}$ used during testing. We observed that applying noise filtering decrease IR RRMSE on average if the input noise intensity during testing were larger than that during training (Figure 7, solid orange and dashed orange lines). Additionally, we observed that updating the empirical estimate of covariance matrix $\hat{R}$ $(\alpha_{R}=0.1>0)$ reduces the IR RRMSE (Figure 7, solid orange and dotted orange lines).
@@
Figure 7: Input reconstruction performance for different input noise intensities. The means and standard deviations among 50 trials are depicted by the lines and shaded regions, respectively.
@@ 4.1 Unsupervised input reconstruction
In Section 2, we demonstrated that input reconstruction (IR) in ESN can be formalized as unsupervised learning (UL), which does not require data from the original inputs.
@@ 4.1.1 ESN information processing ability
This UL-based formalization enhances our understanding of the ESN information processing ability. The ESN readout layer is trained to match the desired ESN output ((2) in Section 2). As the desired output is generally a nonlinear and time-delayed transformation of the input, the input itself can be regarded as the simplest choice for the desired output. In this sense, the IR performance could correspond to the upper limit of the ESN information processing ability. Our finding that IR in ESNs is UL indicates that the upper limit of the ESN information processing ability is independent of the type of input. This natural statement facilitates the independent investigation of two key aspects. First, it raises the question of what kind of ESN properties, such as the characteristics of the activation function $\sigma$ and the parameters $A$ and $B$, are suitable for IR. Second, it highlights the need to understand which properties of the ESN input contribute to the degradation of its raw information processing ability. By addressing these aspects, the UL-based IR formalization provides guidelines for understanding both the universal and task-dependent ESN properties.
@@ 4.1.2 ESN behavior toward the input
We developed an algorithm to compute the readout layer parameter $W_{\mathcal{R}}\in\mathbb{R}^{n_{\text{in}}\times n_{\text{r}}}$, which maps the reservoir state $r_{t}$ to the current input $d_{t}$. Using this, we can numerically obtain the right inverse map of $W_{\mathcal{R}}$ as follows:
$V_{\mathcal{R},\Xi}:=W_{\mathcal{R}}^{+}+(I-W_{\mathcal{R}}^{+}W_{\mathcal{R}})\Xi,$ (14)
where $\Xi\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ is an arbitrary matrix. Suppose that $\text{rank}(W_{\mathcal{R}})=n_{\text{in}}$. Then, by definition, the following equation holds:
$W_{\mathcal{R}}V_{\mathcal{R},\Xi}d_{t}=d_{t}\approx W_{\mathcal{R}}r_{t}.$ (15)
This implies that, for some $\Xi$, $V_{\mathcal{R},\Xi}$ could be a linear approximation of the echo state map (Hart et al., 2020), generalized synchronization (Kocarev and Parlitz, 1996; Lu and Bassett, 2020), or embedding (Sauer et al., 1991; Grigoryeva et al., 2021), which maps $d_{t}$ to $r_{t}$. Thus, our formalization could provide an approach to analyze the ESN behavior toward the input by considering the linear map $V_{\mathcal{R},\Xi}$. A key challenge in this analysis is determining how to calculate the matrix $\Xi$ in (14) so that $V_{\mathcal{R},\Xi}d_{t}\approx r_{t}$ holds. Additionally, it is important to investigate whether this calculation of $\Xi$ can be achieved using UL. Addressing these questions will further clarify the universal behavior of ESNs toward inputs, which should be characterized by inherent properties of ESNs rather than specific inputs.
4.1.3 With uncertain knowledge of ESN parameters
In this study, we assumed complete knowledge of the ESN parameters $A$ and $B$, which leads to UL-based IR. For further analysis, one can consider the generalized case where this knowledge is uncertain. The equation (5) describing the ESN readout layer obtained through UL-based IR indicates that scale and order uncertainty in the ESN input layer results in blind source separation. Specifically, in an ESN, scale and order uncertainty in the input layer directly affects the scale and order uncertainty of the IR output as follows:
$\tilde{W}_{\mathcal{R}}=(BDP)^{+}(\hat{A}-A)=P^{\top}D^{-1}B^{+}(\hat{A}-A)=P^{\top}D^{-1}W_{\mathcal{R}},$ (16)
where $\tilde{W}_{\mathcal{R}}$ represents the ESN readout layer obtained through UL-based IR under uncertainty, and $D$ and $P$ denote a scaling matrix and a permutation matrix, respectively. In other words, the formalization in this study, which incorporates scale and order uncertainty in the ESN input layer, establishes an unsupervised learning algorithm for blind source separation in ESNs. This enables a comparison between our approach and various unsupervised learning algorithms for blind source separation. The references (Choi et al., 2005; Pal et al., 2013; Ansari et al., 2023) provide comprehensive reviews of existing methods. Besides general BSS methods, it is also important to compare our ESN-based approach with previous studies that have implemented BSS using neurocomputational frameworks (Amari and Cichocki, 1998; Isomura and Toyoizumi, 2016; Asabuki and Fukai, 2020; Bahroun et al., 2021; Isomura and Toyoizumi, 2021). We note that our formalization allows for a more general case by introducing further uncertainty in the knowledge of ESN parameters $A$ and $B$. Thus, it is interesting to investigate whether there exist additional types of uncertainty in the knowledge of ESN parameters $A$ and $B$ that still allow for blind source separation in ESNs.
@@ 4.2.1 Standard approach to dynamical system replication
The UL-based formalization encourages us to reconsider the standard approach to dynamical system replication. We observed that the replicated dynamical systems using SL and UL differ because of the numerical errors. These errors are practically unavoidable. For example, numerical errors in the computation of the inverse tanh function are unavoidable because they involve Napier’s constant, which is approximated as a rational number in computers. Similarly, numerical errors in the computation of the Moore-Penrose inverse of matrices are unavoidable because of the singular value decomposition, where singular values smaller than the machine epsilon are typically truncated by default. Thus, we propose treating the replicated dynamical systems obtained from SL and UL as distinct, even though they are theoretically equivalent, based on assumptions made for ESNs, as explained in Section 2. Moreover, we propose considering the replicated dynamical systems obtained from UL as a standard because this approach does not require data for the original ESN input. In contrast, the replicated dynamical systems obtained from SL are treated as special cases that can be implemented only when the original ESN input is exceptionally available. It is worth investigating how the assumption of the original input availability affects characteristics such as dynamical system replication robustness.
4.2.2 Contributions to theoretical neuroscience
@@
A formalization independent of the original ESN input enables the investigation of the computational principles of self-organizing systems, such as the brain. Our perception, sustained by the brain, allows us to understand the external environment through sensory information. All computations required for perception occur within the brain and do not explicitly depend on the true state of the external environment. In this context, our UL-based formalization contributes to further exploration of the computational principles of the brain. In particular, if the activation function $\sigma$ is element-wise and injective, such as the tanh or the identity function, the loss function defined in (6) is equivalent to the following loss function of $\hat{A}$:
@@
$\sum_{t=1}^{T-1}\|\underbrace{r_{t+1}}_{\text{stimulus-evoked activity}}-\underbrace{\sigma(\hat{A}r_{t})}_{\text{internally-predicted activity}}\|^{2}.$ (17)
@@
This reformulation interprets the loss function $L_{\sigma,T}(\hat{A})$ as the discrepancy between the stimulus-evoked activity $r_{t+1}$ and the internally-predicted activity $\sigma(\hat{A}r_{t})$. This perspective parallels prior studies (Rao and Ballard, 1999; Knill and Pouget, 2004; Friston, 2010; Asabuki and Clopath, 2024), which propose that minimizing the discrepancy between these activities reflects a core computational principle underlying perception. We highlight a shared mathematical foundation between the prior studies addressing brain models for perception and the ESNs computational framework. Notably, we deductively derived this loss function from the mathematical observation on IR, hence avoiding heuristic implementation of the loss function (17). Providing that ESN is one of the simplest models capturing the essence of perception based on the loss function (17), we believe the ESN potential as a theoretical model of the brain. To the best of our knowledge, it is remain unclear whether the introduction of additional complexities in the model guarantee qualitative benefits in the tasks treated in this study. We believe that our formulation using ESN provides a foundational basis for future investigations into the detailed properties of both the brain and ESNs.
4.3 Unsupervised foise filtering
@@ 4.3.1 Effective design of ESN with error feedback
$r_{t+1}^{(2\to 1)}$ $=\hat{f}^{(1)}_{\mathcal{R}}(r_{t}^{(2\to 1)})+K(r_{t+1}^{(2)}-\hat{f}^{(1)}_{\mathcal{R}}(r_{t}^{(2\to 1)})),$ (18)
where $K=P(P+R)^{-1}\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ is the stationary Kalman gain, defined using the solution $P$ of the algebraic Riccati equation $P=\hat{A}(P-P(P+R)^{-1}P)\hat{A}^{\top}+Q$. (18) can be reformulated as follows:
@@
$+\underbrace{\left[g(d_{t}^{(2)},r_{t}^{(2)})-g(d_{t}^{(2)},r_{t}^{(2\to 1)})\right]}_{\text{residual}}.$ (19)
@@
We note that (4.3.1) can be viewed as an ESN with IR error feedback (Katori, 2018) defined as follows:
$r_{t+1}:=\underbrace{g(d_{t}^{(2)},r_{t})}_{\text{main dynamics}}+\underbrace{W^{\text{fb}}(d_{t}^{(2)}-W_{\mathcal{D}}r_{t})}_{\text{error feedback}},$ (20)
where $W^{\text{fb}}\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ is the ESN feedback layer. Thus, the proposed noise filtering algorithm induces the design of ESN with IR error feedback. In particular, parameter $W^{\text{fb}}$ in the ESN feedback layer recommended to be set as $(K-I)B$, provided that the linearity assumption of the activation function $\sigma$ is acceptable. This parameter setting, which uses the Kalman gain $K$, implements an intuitively natural strategy: it utilizes information obtained from current input $d_{t}^{(2)}$ only if its noise intensity is smaller than the learning error in the readout layer $W_{\mathcal{R}}$. Therefore, the construction of the filtering algorithm without the explicit use of the ESN input theoretically supports the benefits of error feedback in ESNs, as reported, in (Yonemura and Katori, 2021; Sato et al., 2024), for example.

### v4

@@ 3.2.3 Numerical Experiments
We generated inputs $d_{t}^{(1)}=d_{t}+\epsilon_{t}^{(1)}$ and $d_{t}^{(2)}=d_{t}+\epsilon_{t}^{(2)}$, where $\epsilon_{t}^{(1)}$ and $\epsilon_{t}^{(2)}$ are zero-mean white Gaussian noise with variance $\text{tr}(\Sigma^{(1)})=0.01$ and $\text{tr}(\Sigma^{(2)})$ is varied over 10 logarithmically spaced values between $10^{-2}$ and $10^{1}$. We used a 30-dimensional ESN with the element-wise $\tanh$ as the activation function $\sigma.$ We set the elements of the input layer parameter $A\in\mathbb{R}^{30\times 1}$ using i.i.d samples obtained from $\mathcal{N}(0,0.02)$. Let $B_{0}\in\mathbb{R}^{30\times 30}$ be a matrix with elements being i.i.d samples from $\mathcal{N}(0,1)$. We set $B=0.9\cdot\rho(B_{0})\cdot B_{0}$, where $\rho(B_{0})$ is the spectral radius of $B_{0}$. For each ESN implementation and noisy inputs $\mathcal{D}_{1,5000}^{(1)}$ and $\mathcal{D}_{1,3000}^{(2)}$, we calculated reservoir states $\mathcal{R}_{1,5001}^{(1)}$ and $\mathcal{R}_{1,3001}^{(2)}$ with the zero vector as the initial reservoir state. We applied Algorithm 2 to obtain $W_{R}^{(1)},\hat{f}_{\mathcal{R}}^{(1)}$ and $\hat{Q}$ from $\mathcal{R}_{1,5001}^{(1)}$. Then, we applied the adaptive EnKF Algorithm 3 with $\alpha_{R}=0.01$ and $M=300$ to obtain filtered reservoir states $r_{t}^{(2\to 1)}$ from the sequence of contaminated reservoir states $r_{t}^{(2)}$. We conducted simulations for 10 different implementations of the ESN parameters $A$ and $B$ and noisy inputs $\mathcal{D}_{1,5000}^{(1)}$ and $\mathcal{D}_{1,3000}^{(2)}$. We calculated the relative root mean squared error (RRMSE) to evaluate the IR performance.
We additionally evaluated the proposed method under non-Gaussian observation noise, including heavy-tailed distributions such as the Student’s $t$ and Cauchy distributions. Results demonstrating consistent performance are reported in the Supplementary Material.
@@
Figure 6 shows a typical numerical example of the ESN inputs and outputs using IR with noise filtering. We used low-intensity (Figure 6(a)) and high-intensity (Figure 6(b)) noise inputs for training and testing, respectively. In this trial, the RRMSE of IR was reduced from 1.031.031.03 to 0.500.500.50 by applying noise filtering (Figure 6(c) and (d)).
@@
Figure 6: Typical example of input reconstruction with noise filtering. (a) ESN input $d_{t}^{(1)}$ with low-intensity noise $(\text{tr}(\Sigma^{(1)})=0.01)$ used in prior training. (b) ESN input $d_{t}^{(2)}$ with high-intensity noise $(\text{tr}(\Sigma^{(2)})=1)$ used in testing. (c) ESN output $W_{\mathcal{R}}^{(1)}r_{t}^{(2)}$ before filtering. (d) ESN output $W_{\mathcal{R}}^{(1)}r_{t}^{(2\to 1)}$ after filtering. The original input $d_{t}$ is shown as a reference (black dashed line) in all panels.
@@
Figure 7 shows the average behavior of the IR RRMSE across 10 trials for each noise intensity of inputs $d_{t}^{(2)}$ used during testing. We observed that applying noise filtering decreases IR RRMSE on average when the input noise intensity during testing is higher than that during training (Figure 7, solid orange and dashed orange lines). We also found that adaptively updating the covariance estimate $\hat{R}$ (with $\alpha_{R}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.01}$) significantly lowers the RRMSE of input reconstruction when the test input is contaminated with low-intensity noise ($\sqrt{\text{tr}(\Sigma^{(2)})}<1$). Noise filtering without adaptation of covariance matrix $\hat{R}_{t}$ also works when the input noise intensity during testing $\sqrt{\text{tr}\Sigma^{(2)}}<1$ is $10^{0}$ to $10^{1}$ due to the suitability of the initial covariance matrix estimate $\hat{R}_{0}$ to the noise level in the test data.
@@
Figure 7: Input reconstruction performance for different input noise intensities. The means and standard deviations among 10 trials are depicted by the lines and shaded regions, respectively.
@@ 4.1 Unsupervised Input Reconstruction
In Section 2, we demonstrated that IR in ESN can be formalized as UL, which does not require data from the original inputs.
@@ 4.1.1 ESN information processing ability
This UL-based formalization enhances our understanding of the ESN information processing ability. The ESN readout layer is trained to match the desired ESN output (see Equation 2.2). As the desired output is generally a nonlinear and time-delayed transformation of the input, the input itself can be regarded as the simplest choice for the desired output. In this sense, the IR performance could correspond to the upper limit of the ESN information processing ability. Our finding that IR in ESNs is UL indicates that the upper limit of the ESN information processing ability is independent of the type of input. This perspective allows us to independently investigate the following questions:
  • What kind of ESN properties, regarding the activation function $\sigma$ and parameters $A$ and $B$, are suitable for IR?
  • Which properties of the ESN input degrade its raw information processing ability?
Thus, the UL-based IR formalization provides guidelines for understanding both the universal and task-dependent ESN properties.
@@ 4.1.2 ESN behavior toward the input
We developed an algorithm to compute the readout layer parameter $W_{\mathcal{R}}\in\mathbb{R}^{n_{\text{in}}\times n_{\text{r}}}$, which maps the reservoir state $r_{t}$ to the current input $d_{t}$. Using this, we can numerically construct a parametric family of candidate inverse maps of $W_{\mathcal{R}}$ as follows:
$V_{\mathcal{R},\Xi}:=W_{\mathcal{R}}^{+}+(I-W_{\mathcal{R}}^{+}W_{\mathcal{R}})\Xi,$ (4.1)
where $\Xi\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ is an arbitrary matrix. Suppose that $\operatorname{rank}(W_{\mathcal{R}})=n_{\text{in}}$, i.e., $W_{\mathcal{R}}$ is full row rank. Then the Moore–Penrose inverse satisfies $W_{\mathcal{R}}W_{\mathcal{R}}^{+}=I_{n_{\text{in}}}$, and since $W_{\mathcal{R}}(I-W_{\mathcal{R}}^{+}W_{\mathcal{R}})=0$, we obtain $W_{\mathcal{R}}V_{\mathcal{R},\Xi}=I_{n_{\text{in}}}$. That is, $V_{\mathcal{R},\Xi}$ is a right inverse of $W_{\mathcal{R}}$. Furthermore, since the readout layer $W_{\mathcal{R}}$ is trained to minimize the reconstruction error of the input, we have $d_{t}\approx W_{\mathcal{R}}r_{t}$ by construction. Under this assumption and the full row rank condition, it is plausible that $r_{t}\approx V_{\mathcal{R},\Xi}d_{t}$ for some choice of $\Xi$. This suggests that $V_{\mathcal{R},\Xi}$ may serve as a linear surrogate for the inverse mapping from input to reservoir state. From this perspective, $V_{\mathcal{R},\Xi}$ can be interpreted as a linear approximation of the echo state map (Hart et al.,, 2020), generalized synchronization (Kocarev and Parlitz,, 1996; Lu and Bassett,, 2020), and embedding (Sauer et al.,, 1991; Grigoryeva et al.,, 2021), all of which relate the input signal $d_{t}$ to the internal reservoir state $r_{t}$. This connection provides a principled framework for analyzing how ESNs encode input signals into internal states through approximate inverse mappings. In future work, the following issues need to be addressed:
  • How can we calculate matrix $\Xi$ in Equation 4.1 to satisfy $V_{\mathcal{R},\Xi}d_{t}\approx r_{t}$?
  • Is the calculation of $\Xi$ possible using UL?
@@ 4.2.1 Standard approach to dynamical system replication
The UL-based formalization encourages us to reconsider the standard approach to dynamical system replication. We observed that the replicated dynamical systems using SL and UL differ because of numerical errors. These errors are practically unavoidable. For example, numerical errors in the computation of the inverse tanh function are unavoidable because they involve Napier’s constant, which is approximated as a rational number in computers. Similarly, numerical errors in the computation of the Moore-Penrose inverse of matrices are unavoidable because of the singular value decomposition, where singular values smaller than the machine epsilon are typically truncated by default. Thus, we propose treating the replicated dynamical systems obtained from SL and UL as distinct, even though they are theoretically equivalent, based on assumptions made for ESNs, as explained in Section 2. Moreover, we propose considering the replicated dynamical systems obtained from UL as a standard because this approach does not require data for the original ESN input. In contrast, the replicated dynamical systems obtained from SL are treated as special cases that can be implemented only when the original ESN input is exceptionally available. It is worth investigating how the assumption of the original input availability affects characteristics such as dynamical system replication robustness.
4.2.2 Relations to theoretical neuroscience
@@
A formalization independent of the original ESN input enables the investigation of the computational principles of self-organizing systems, such as the brain. Our perception, sustained by the brain, allows us to understand the external environment through sensory information. All computations required for perception occur within the brain and do not explicitly depend on the true state of the external environment. In this context, our UL-based formalization contributes to further exploration of the computational principles of the brain. In particular, if the activation function $\sigma$ is element-wise and injective, such as the tanh or the identity function, the loss function defined in Equation 2.6 is equivalent to the following loss function of $\hat{B}$:
@@
$\sum_{t=1}^{T-1}\|\underbrace{r_{t+1}}_{\text{stimulus-evoked activity}}-\underbrace{\sigma(\hat{B}r_{t})}_{\text{internally-predicted activity}}\|^{2}.$ (4.2)
@@
This reformulation interprets the loss function $L_{\sigma,T}(\hat{B})$ as the discrepancy between the stimulus-evoked activity $r_{t+1}$ and the internally-predicted activity $\sigma(\hat{B}r_{t})$. This perspective parallels prior studies (Rao and Ballard,, 1999; Knill and Pouget,, 2004; Friston,, 2010; Asabuki and Clopath,, 2024), which propose that minimizing the discrepancy between these activities reflects a core computational principle underlying perception. We highlight a shared mathematical foundation between the prior studies addressing brain models for perception and the ESN’s computational framework. Notably, we deductively derived this loss function from the mathematical observation on IR, hence avoiding heuristic implementation of the loss function 4.2. Provided that ESN is one of the simplest models capturing the essence of perception based on the loss function 4.2, we believe the ESN potential as a theoretical model of the brain. To the best of our knowledge, it remains unclear whether the introduction of additional complexities in the model guarantee qualitative benefits in the tasks treated in this study. We believe that our formulation using ESN provides a foundational basis for future investigations into the detailed properties of both the brain and ESNs.
4.3 Unsupervised Noise Filtering
@@ 4.3.1 Effective design of ESN with error feedback
$r_{t+1}^{(2\to 1)}$ $=\hat{f}^{(1)}_{\mathcal{R}}(r_{t}^{(2\to 1)})+K(r_{t+1}^{(2)}-\hat{f}^{(1)}_{\mathcal{R}}(r_{t}^{(2\to 1)})),$ (4.3)
where $K=P(P+R)^{-1}\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ is the stationary Kalman gain, defined using the solution $P$ of the algebraic Riccati equation $P=\hat{B}(P-P(P+R)^{-1}P)\hat{B}^{\top}+Q$. Equation 4.3 is reformulated as follows:
@@
$+\underbrace{\left[g(d_{t}^{(2)},r_{t}^{(2)})-g(d_{t}^{(2)},r_{t}^{(2\to 1)})\right]}_{\text{residual}}.$ (4.4)
@@
We note that Equation 4.3.1 can be viewed as an ESN with IR error feedback (Katori,, 2018) defined as follows:
$r_{t+1}:=\underbrace{g(d_{t}^{(2)},r_{t})}_{\text{main dynamics}}+\underbrace{W^{\text{fb}}(d_{t}^{(2)}-W_{\mathcal{D}}r_{t})}_{\text{error feedback}},$ (4.5)
[…]
4.3.2 Extensions toward Non-stationary Noise
In filtering problems, the system noise covariance matrix $Q$ and the observation noise covariance matrix $R$ are typically incorporated into the algorithm to reflect the probabilistic characteristics of the system state and observations.
In the state-space model introduced in this study (Equation 3.3), system noise is formulated as the modeling error of a dynamical system that has been pretrained in an unsupervised manner using an ESN. Accordingly, the matrix $Q$ is fixed once pretraining is completed and does not vary with the noise characteristics of the input sequence. In contrast, observation noise is defined as the discrepancy between the actual ESN state and its estimate at each time step, and thus the matrix $R$ is influenced by the noise properties of the ESN input and cannot be determined a priori.
This study focuses on a stationary setting in which the true observation noise covariance is unknown but constant. Under this assumption, we numerically verified the effectiveness of adaptive filtering for online estimation of $R$ (Figure 7, solid orange and dashed orange lines). More generally, input noise may be non-stationary. The adaptive filter (Mehra,, 1972) employed in this work is one standard approach to address such cases. The observed improvement from adaptation (Figure 7, solid orange and dotted black lines) suggests that the proposed algorithm is capable of estimating unknown and potentially time-varying covariance matrices, and therefore remains effective as long as the temporal variation of noise is sufficiently slow.
Alternative approaches to non-stationary noise filtering include the Switching Kalman Filter (Murphy,, 1998), which simultaneously runs multiple filters based on different state-space models, and Augmented State Kalman Filtering (Deaves,, 1999), which assumes prior knowledge of the covariance dynamics and incorporates them as part of the system state. Extending our framework to incorporate such methods may be a fruitful direction for future development, depending on the characteristics of the target application.
4.4 Limitations
4.4.1 Theoretical requirements for enabling UL formulation of IR in ESNs
This study has clarified the theoretical assumptions on ESN parameters required to enable UL formulation of IR within ESNs. These assumptions fall into two main categories: (1) regularity conditions on the ESN and (2) full accessibility of the internal structure of the ESN.
The first requirement concerns the full column rank dondition of the input layer matrix $A$ and the invertibility of activation function $\sigma$, which enables the reconstruction of the input time series $\mathcal{D}_{1,T}$ from the internal ESN state $\mathcal{R}_{1,T+1}$. While these conditions may seem restrictive in theory, they are often satisfied in practice when using randomly generated input matrices and standard activation functions such as the identity or hyperbolic tangent function, both of which are commonly employed in ESN implementations. We numerically confirmed that the non-invertibility of activation functions degrades the performance of IR, with ReLU-based ESNs achieving an RRMSE of approximately $10^{-2}$ (Figure S3). Although this is worse than the near-perfect performance under invertible functions, it remains significantly better than chance level (RRMSE = 1). Accordingly, we suggest that nonlinear invertible activation functions, such as the hyperbolic tangent, are more tractable than non-invertible ones under the proposed theory of IR in ESN.
The second requirement is full knowledge of the internal components of the ESN, specifically $A$, $B$, and $\sigma$. This assumption allows us to reformulate IR as a UL problem by substituting the need for target outputs with knowledge of the model parameters. The practicality of this assumption depends on the specific application scenario. As highlighted in this study, digital reservoirs such as ESNs typically provide full access to internal structures, including connectivity matrices and activation functions. Our findings support the utility of leveraging such prior information, as they demonstrate that IR, which is typically formulated as an SL task, can instead be reformulated as a UL problem when the internal components are known. In contrast, physical reservoir systems (Tanaka et al.,, 2019), including analog circuits (Appeltant et al.,, 2011), soft robotic systems (Nakajima et al.,, 2015), and chemical reaction systems (Baltussen et al.,, 2024), often operate under practical constraints that limit precise design or access to internal dynamics. Historically, the reservoir computing paradigm has implicitly treated the reservoir as a black-box component due to its fixed and untrained nature (Jaeger,, 2001; Maass et al.,, 2002). As a result, explicit modeling of the reservoir’s internal dynamics has received little attention. However, this perspective tends to conflate two distinct assumptions: that the reservoir is untrained, and that its internal structure is inaccessible. While the former remains a central tenet of reservoir computing, the latter is not inherently necessary. In this context, the present work provides a new perspective, pointing toward research directions that take advantage of the internal behavior of physical reservoirs. While specifying exact connectivity matrices or activation functions in such systems may be challenging, reproducibility of internal dynamics across fabricated instances may still be attainable. For instance, analog circuits are typically engineered to behave deterministically, apart from variability due to manufacturing tolerances or thermal noise. Soft robots exhibit structural consistency in their dynamics once fabricated, and chemical reaction systems are constrained by physical principles such as conservation of mass, which impose qualitative regularities. The central question is whether one can construct a reliable dynamical model that captures these deterministic features. Given recent advances in the simulation of physical systems, we believe that our UL-based formulation has the potential to be extended to the domain of physical reservoir computing.
An even more challenging setting arises when the identity and structure of the reservoir system are entirely unknown. In such cases, our formulation introduces a novel system identification problem: inferring the reservoir model solely from observed reservoir state trajectories. Although we do not propose a specific solution strategy in this study, several theoretical challenges are evident. Assuming that the target reservoir can be approximated by a suitable ESN, the problem becomes one of estimating the activation function and weight matrices from the observed reservoir states. This raises fundamental issues such as identifiability, for example, scale or permutation indeterminacy in the input weight matrix when the input time series is unobserved. It also introduces an infinite-dimensional estimation problem due to the unknown activation function, placing the task within the domain of semi-parametric inference (Li and Duan,, 1989; Bickel et al.,, 1993). Determining which classes of ESNs are identifiable from reservoir state sequences alone, and whether such classes retain expressive capabilities such as universal approximation, represent promising avenues for future research suggested by our formulation.
4.4.2 Performance evaluations
This study does not include a performance comparison between our proposed method and existing algorithms for IR, such as those developed in the context of blind source separation (e.g., Bell and Sejnowski,, 1995; Amari and Cichocki,, 1998; Isomura and Toyoizumi,, 2016; Asabuki and Fukai,, 2020; Bahroun et al.,, 2021), dynamical system replication using machine learning models other than RC (e.g., Chen et al.,, 2018; Dueben and Bauer,, 2018; Scher and Messori,, 2019), or model-free noise filtering (e.g., Vincent et al.,, 2008; Hamilton et al.,, 2016). This omission is intentional: our primary aim is not to develop a new state-of-the-art algorithm, but to establish a theoretical framework for UL-based IR and its connections to related tasks. Our contribution lies in identifying the conditions under which IR can be achieved without supervision in ESNs, and in demonstrating that related tasks such as dynamical system replication and noise filtering can be reformulated within this unified framework. These reformulations enable the transfer of existing theoretical guarantees into the UL setting.
Nonetheless, we acknowledge that empirical evaluation on real-world tasks is essential for assessing practical utility. Future work should explore the performance and limitations of UL-based IR, particularly under noisy or data-constrained conditions.